OpenStudy (anonymous):

Andre received $60 for his birthday and put it in his piggy bank. Each week, he puts 3 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 3w + 60 where w is the number of weeks he has saved 3 dollars. His sister Lana makes a table of reasonable values for w. Lana selects 12 as a value for w. What is the corresponding table entry for D? A. 372 B. 96 C. 75 D. 64

OpenStudy (anonymous):

@cj49

OpenStudy (anonymous):

The answers are different so its not 186.

OpenStudy (cj49):

3(12)+60=36+60=96